42. use half angle formula to find cos sin and tan of 67 degrees and 30'
Answer by stanbon(75887)

use half angle formula to find cos sin and tan of 67 degrees and 30'
-------------
sin(67 30') = sin(135/2)
-------
Formula: sin(135/2) = sqrt[(1-cos(135))/2]
= sqrt[(1-(1/sqrt(2))/2] = sqrt[0.85355..]= 0.92388
====================================================
cos(67 30') = cos(135/2)
-------
Formula: cos(135/2) = sqrt[(1+cos(135))/2]
= sqrt[0.146447] = 0.38268..
====================================================
tan(67 30') = tan(135/2)
------
Formula: tan(135/2) = (1-cos(135))/2 = 0.85355...
====================================================
Cheers,
Stan H.